OpenStudy (anonymous):

Eight less than a number doubled is twenty more than the number. What is the number?

OpenStudy (evanhelps):

2x-8=20+x 2x=28+x x=28

OpenStudy (evanhelps):

You just have to set up the equation and then solve it

OpenStudy (anonymous):

thanks